The sum of two numbers is 66. The larger number is three less than twice the
smaller number. Find the numbers.

Answer by deborabr(181)

x+y=66
2y-3=x --> substituting x --> (2y-3)
(2y-3)+y=66
2y-3+y=66
2y+y=66+3
3y=69
y=69/3
y=23
replacing the y to find the x:
2y-3=x
2.23-3=x
46-3=x
43=x
Answer: 23 e 43